# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v1

@@ 3 RBK without Preprocessing
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (14)
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq\left(1-\alpha\right)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}\\ +\frac{1}{\alpha^{2}(T-T_{b})}\;\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$ (15)
To our knowledge, this is the first result for inconsistent linear systems that characterizes the exact solution to which the randomized block Kaczmarz iterates converge, albeit in a Monte Carlo sense. The $O(1/T)$ convergence rate for the tail-averaged bound is optimal for row-access methods, and a reasonable default for the burn-in time is $T_{b}=T/2$; see (Epperly et al., 2024) for a more thorough discussion of these points. The proof of Theorem 3.1, which takes advantage of an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$, can be found in Appendix B.
Unfortunately, Theorem 3.1 does not imply robust convergence for general problems. For one, the convergence parameter $\alpha$, which affects every term in the bound, is difficult to analyze in general. Worse, for problems containing nearly singular blocks $A_{S}$, the weight matrix $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(A_{S}A_{S}^{\top})^{+}I_{S}\right]$ can be arbitrarily poorly conditioned and the variance term $\mathbb{E}\,_{S\sim\rho}||A_{S}^{+}r^{(\rho)}_{S}||^{2}$ can be arbitrarily large. See Figure 2 for numerical examples that manifest these problems.
In the next subsection, we show that these problems can be overcome when the data arises from a Gaussian distribution. A more general solution is provided by the ReBlocK algorithm introduced in Section 4.
@@ 3.1 Linear Least-squares with Gaussian Data
$\begin{bmatrix}a_{i}^{\top}&b_{i}\end{bmatrix}\sim\mathcal{N}(0,Q)$ (16)
@@
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$ (17)
@@
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the randomly generated finite problem (16). Then the results of Theorem 3.1 hold with
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@
$\lim_{m\rightarrow\infty}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}\leq\frac{200}{\sigma_{2k}^{2}}\cdot\lim_{m\rightarrow\infty}\frac{\left\lVert r\right\rVert^{2}}{m}.$ (18)
Finally, the convergence parameter $\alpha$ satisfies
$\lim_{m\rightarrow\infty}\alpha\geq C_{n,k}\max\left\{\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}},\max_{2\leq\ell<k}\frac{(\ell-1)\sigma_{n}^{2}}{\sum_{i\geq k-\ell-1}\sigma_{i}^{2}}\right\}$ (19)
with $C_{n,k}\rightarrow 1$ as $n\rightarrow\infty$ for fixed $k$.
Thus, in the case of Gaussian data and large $m$, the ordinary least-squares solution is recovered, the variance of the iterates is bounded, and the convergence rate $\alpha$ improves at least linearly with the block size $k$. In addition, the following corollary, which is based on Corollary 3.4 of (Dereziński & Rebrova, 2024), shows that RBK converges much faster than mSGD for polynomially decaying singular values.
@@
$\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq Ck^{\beta}\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}},\;\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq k\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}}$ (20)
@@
Similarly, the faster convergence rate of RBK over mSGD extends to exponentially decaying singular values. The proofs of Theorems 3.2 and 3.3, provided in Appendix B, rely on a connection between RBK-U for Gaussian data and sketch-and-project with Gaussian sketch matrices. These results generalize the techniques of (Derezinski & Mahoney, 2021; Dereziński & Rebrova, 2024) to the case of inconsistent linear systems, improving upon the results of (Rebrova & Needell, 2021) in terms of both the convergence rate and the variance. It is worth noting that Gaussian data is just one way to generate a benign or “incoherent” problem. We expect that Theorems 3.2 and 3.3 could be generalized to other cases such as problems that have been preprocessed using a randomized Hadamard transform, potentially extending the results of (Dereziński & Yang, 2024) to the inconsistent case.
@@ 3.3 Numerical Demonstration
We test our theoretical results for the RBK-U and mSGD algorithms on two problems with Gaussian data, with results in Figure 1. We apply tail averaging to each algorithm to observe convergence beyond the variance horizon. As expected, tail-averaged RBK-U (TA-RBK-U) converges much more rapidly than tail-averaged mSGD (TA-mSGD) in the presence of fast singular value decay. Further details on these experiments and a link to the code can found in Appendix F.
@@ 4 Robustness through Regularization
To address the shortcomings of RBK-U in the case of general data, we propose to incorporate regularization into the RBK algorithm. A natural way to do this is to replace the RBK iteration with a stochastic proximal point iteration (Asi & Duchi, 2019), namely
$x_{t+1}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left[\left\lVert A_{S_{t}}x-b_{S_{t}}\right\rVert^{2}+\lambda k\left\lVert x-x_{t}\right\rVert^{2}\right].$ (21)
@@
$x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}),$ (22)
and that the RBK iteration (5) is recovered in the limit $\lambda\rightarrow 0$. To incorporate a mild regularization without significantly slowing down the convergence, we propose to use a small, constant value of $\lambda$ throughout the algorithm. We suggest $\lambda=0.001$ as a practical default value, but to obtain optimal performance the value may need to be tuned on a case-by-case basis; see Section F.1.1 for further discussion. We refer to the resulting scheme as the regularized block Kaczmarz method, or ReBlocK.
@@
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (23)
@@
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\frac{\left\lVert r^{(\rho)}\right\rVert^{2}}{m}.$ (24)
@@
$\left\lVert r^{(\rho)}\right\rVert^{2}\leq\left(1+\frac{N}{\lambda}\right)\left\lVert r\right\rVert^{2}.$ (25)
Note that the values of $x^{(\rho)}$ and $\alpha$ here differ from those in Section 2 due to the different choice of $M(A_{S})$ used by ReBlocK. To our knowledge, Theorem 4.1 is the first result characterizing the convergence of a stochastic proximal point algorithm with a constant step size for inconsistent linear systems. The advantage relative to RBK-U is that ReBlocK-U is able to control both the variance of the iterates and the conditioning of the weight matrix $\overline{W}$ in terms of the reciprocal of the regularization parameter $\lambda$. As a result, the algorithm converges robustly even when the problem contains many nearly singular blocks $A_{S}$. It is worth noting that these advantages could also be attained by truncating the singular values that are smaller than $\lambda$ in the RBK iteration (5). However, the ReBlocK iteration is additionally justified by the fact that it is cheaper to implement than the RBK iteration; see Section 4.1 for further discussion of this point.
The proof of Theorem 4.1 is provided in Appendix C. Relative to the proof of Theorem 3.1, the proof of Theorem 4.1 is more complicated because the ReBlocK iteration does not lead to an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$. Instead, the proof relies on a bias-variance decomposition inspired by (Défossez & Bach, 2015; Jain et al., 2018b; Epperly et al., 2024), which also leads to the extra factors of $2$ in the convergence bound.
We are not yet able to analyze the convergence rate of ReBlocK-U, even in the case of Gaussian data, as to our knowledge there is no existing work bounding the quality of a regularized Gaussian sketch. However, convergence to the ordinary least-squares solution is obtained for a broader class of noisy linear least-squares problems; see Appendix D. Additionally, a fast rate of convergence to the ordinary least-squares solution can be shown when sampling from an appropriate determinantal point process; see Appendix E.
@@ 4.1 Implementation Details
To implement the ReBlocK iterations (7), we directly calculate the $k\times k$ matrix $A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI$ and then use a Cholesky-based linear system solver to calculate $(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x)$. This computation is stable as long as $\lambda$ is not chosen to be too small. Using this approach, the most expensive part of the ReBlocK iteration is calculating $A_{S_{t}}A_{S_{t}}^{\top}$, which has an asymptotic cost of $O(nk^{2})$ just like RBK. Nonetheless, in practice the matrix-matrix multiplication for ReBlocK can be significantly cheaper than the SVD used for RBK. For example, in the experiments of Section 5, ReBlocK iterations are about five times faster than RBK iterations, as reported in Section F.2.1. For the largest problems, the ReBlocK iterations could be further accelerated using iterative solvers; see for example Section 4.1 of (Dereziński et al., 2025).
@@ 4.2 Numerical Demonstration
We now test the ReBlocK-U algorithm on two problems whose columns are discretized representations of continuous functions, leading the matrices $A$ to contain many nearly singular blocks of rows. We set $\lambda=0.001$ and generate the inconsistency using random Gaussian noise. We find that TA-RBK-U is unstable for these problems, as expected. Furthermore, TA-ReBlocK-U converges much faster than TA-mSGD in the presence of rapid singular value decay. Further details on these experiments can be found in Section F.1, and an exploration of the effect of choosing different values of $\lambda$ can be found in Section F.1.1.
@@
Figure 3: Comparison of methods for calculating natural gradient directions for a small neural network. The network parameters $\theta$ are taken from three snapshots of a single training run, with one snapshot from the “pre-descent” phase before the loss begins to decrease (Left), one snapshot from the “descent” phase during which the loss decreases rapidly (Middle), and one snapshot from the “post-descent” phase when the loss has stopped decreasing significantly (Right). The algorithms are measured in terms of their progress towards reducing the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$ for the least-squares problem (31), which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. The burn-in time is set to $T_{b}\approx T/20$ in each case, as indicated by the vertical dotted line.
@@ 5 Natural Gradient Optimization
Our original motivation for this work comes from the problem of training deep neural networks using natural gradient descent (Amari, 1998), which is based on an efficient natural gradient induced by a problem-dependent Riemannian metric. Natural gradient descent has been studied extensively in the machine learning community; see (Martens & Grosse, 2015; Ren & Goldfarb, 2019; Martens, 2020). Furthermore, there is increasing evidence that natural gradient methods can improve the accuracy when training neural networks to solve physical equations. See (Müller & Zeinhofer, 2023; Dangel et al., 2024) for applications to physics-informed neural networks and (Pfau et al., 2020; Schätzle et al., 2023) for applications to neural network wavefunctions.
@@
$\min_{\theta}L(\theta),\quad L(\theta)\coloneqq\frac{1}{2}\int_{\Omega}(f_{\theta}(s)-f(s))^{2}{\rm d}s,$ (26)
@@
$\theta\leftarrow\theta-\eta G_{N},\quad G_{N}\coloneqq F^{-1}\nabla_{\theta}L(\theta),$ (27)
@@
$F=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)\nabla_{\theta}f_{\theta}(s)^{\top}{\rm d}s=J^{\top}J,$ (28)
@@
$\nabla_{\theta}L(\theta)=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)(f_{\theta}(s)-f(s)){\rm d}s=J^{\top}[f_{\theta}-f].$ (29)
@@
Calculating the natural gradient direction $G_{N}$ using (27) requires a linear solve against the $n\times n$ matrix $F$, which is very challenging in realistic settings when $n\geq 10^{6}.$ This has motivated the development of approximate schemes such as (Martens & Grosse, 2015). An alternative approach is to reformulate $G_{N}$ using the structure of $F$ and $\nabla_{\theta}L(\theta)$:
$G_{N}$ $=(J^{\top}J)^{-1}J^{\top}[f_{\theta}-f]$ (30)
$=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Jx-[f_{\theta}-f]\right\rVert^{2},$ (31)
where the norm in the final expression is the $L_{2}$-norm in function space. This least-squares formulation has been pointed out for example by (Martens, 2020; Chen & Heyl, 2024; Goldshlager et al., 2024), with the work of Chen and Heyl empowering major advances in the field of neural quantum states. A major goal of the current work is to provide a more solid foundation for the development of natural gradient approximations along these lines.
The natural way to access the data when training a neural network is to sample a set of points in the domain $\Omega$ and evaluate the target function $f$, the network outputs $f_{\theta}$, and the network gradients $\nabla_{\theta}f_{\theta}$ at the sampled points. This is precisely equivalent to sampling a small subset of the rows of (31), which motivates the consideration of row-access least-squares solvers for calculating natural gradient directions. Furthermore, both empirical evidence from scientific applications (Park & Kastoryano, 2020; Wang et al., 2022) and theoretical evidence from the literature on neural tangent kernels (Bietti & Mairal, 2019; Ronen et al., 2019; Cao et al., 2019) suggest that $J$ should be expected to exhibit fast singular value decay, motivating the possibility that RBK and ReBlocK could converge rapidly when solving (31). Note that these observations translate straightforwardly from the simple function learning problem (26) to the realistic problems of training physics-informed neural neteworks or neural network wavefunctions.

### v2

@@ 3 RBK without Preprocessing
(14) $\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$
@@
(15) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq\left(1-\alpha\right)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{\alpha^{2}(T-T_{b})}\;\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$
To our knowledge, this is the first result for inconsistent linear systems that characterizes the exact solution to which the randomized block Kaczmarz iterates converge, albeit in a Monte Carlo sense. The $O(1/T)$ convergence rate for the tail-averaged bound is optimal for row-access methods, and a reasonable default for the burn-in time is $T_{b}=T/2$; see [18] for a more thorough discussion of these points. The proof of Theorem 3.1, which takes advantage of an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$, can be found in Appendix B.
Unfortunately, Theorem 3.1 does not imply robust convergence for general problems. For one, the convergence parameter $\alpha$, which affects every term in the bound, is difficult to analyze in general. Worse, for problems containing nearly singular blocks $A_{S}$, the weight matrix $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(A_{S}A_{S}^{\top})^{+}I_{S}\right]$ can be arbitrarily poorly conditioned and the variance term $\mathbb{E}\,_{S\sim\rho}||A_{S}^{+}r^{(\rho)}_{S}||^{2}$ can be arbitrarily large. See Fig. 3 for numerical examples that manifest these problems.
In the next subsection, we show that these problems are avoided when the data arises from a Gaussian distribution. A more general solution is provided by the ReBlocK algorithm introduced in Section 4.
@@ 3.1 Linear Least-squares with Gaussian Data
(16) $\begin{bmatrix}a_{i}^{\top}&b_{i}\end{bmatrix}\sim\mathcal{N}(0,Q)$
@@
(17) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$
@@
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the randomly generated finite problem (16), where $m$ denotes the total number of generated samples and $k$ is the block size. Then the results of Theorem 3.1 hold with
(18) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
In addition, the convergence parameter $\alpha$ satisfies
(19) $\lim_{m\rightarrow\infty}\alpha\geq C_{n,k}\max\left\{\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}},\max_{2\leq\ell<k}\frac{(\ell-1)\sigma_{n}^{2}}{\sum_{i\geq k-\ell-1}\sigma_{i}^{2}}\right\}$
with $C_{n,k}\rightarrow 1$ as $n\rightarrow\infty$ for fixed $k$.
@@
(20) $\lim_{m\rightarrow\infty}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}\leq\frac{200}{\sigma_{2k}^{2}}\cdot\lim_{m\rightarrow\infty}\frac{\left\lVert r\right\rVert^{2}}{m}.$
In the above theorem, the randomness in the Gaussian data is handled by first assuming a fixed number of samples $m$, and then considering the asymptotic limit $m\rightarrow\infty$. The result is that in the case of Gaussian data, the ordinary least-squares solution is recovered, the convergence rate $\alpha$ improves at least linearly with the block size $k$, and the variance of the iterates is bounded. In addition, the following corollary, which is based on Corollary 3.4 of [14], shows that RBK converges much faster than mSGD for polynomially decaying singular values.
@@
(21) $\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq Ck^{\beta}\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}},\;\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq k\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}}$
@@
Similarly, the faster convergence rate of RBK over mSGD extends to exponentially decaying singular values. The proofs of Theorems 3.2 and 3.3, provided in Appendix B, rely on a connection between RBK-U for Gaussian data and sketch-and-project with Gaussian sketch matrices. These results generalize the techniques of [13, 14] to the case of inconsistent linear systems, improving upon the results of [37] in terms of both the convergence rate and the variance. It is worth noting that Gaussian data is just one way to generate a benign or “incoherent” problem. We expect that Theorems 3.2 and 3.3 could be generalized to other cases such as problems that have been preprocessed using a randomized Hadamard transform, potentially extending the results of [15] to the inconsistent case.
@@ 3.3 Numerical Demonstration
We verify our theoretical results for the RBK-U and mSGD algorithms on two problems with Gaussian data, with results in Fig. 1. We apply tail averaging to each algorithm to observe convergence beyond the variance horizon. As expected, tail-averaged RBK-U (TA-RBK-U) converges much more rapidly than tail-averaged mSGD (TA-mSGD) in the presence of fast singular value decay. We additionally explore the effect of choosing different burn-in periods for TA-RBK-U, with results in Fig. 2. As expected, optimal performance is attained by setting $T_{b}$ to be just slightly after the iterates reach their fixed convergence horizon. The code for all of our experiments is linked in the supplementary material, and further details on these particular experiments can found in F.1.
@@
Figure 2: Behavior of TA-RBK-U with different burn-in periods $T_{b}$. The problem and the reported error quantity are the same as in the left panel of Fig. 1.
@@ 4 Robustness through Regularization
To address the shortcomings of RBK-U in the case of general data, we propose to incorporate regularization into the RBK algorithm. A natural way to do this is to replace the RBK iteration with a stochastic proximal point iteration [2], namely
(22) $x_{t+1}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left[\left\lVert A_{S_{t}}x-b_{S_{t}}\right\rVert^{2}+\lambda k\left\lVert x-x_{t}\right\rVert^{2}\right].$
@@
(23) $x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}),$
and that the RBK iteration (5) is recovered in the limit $\lambda\rightarrow 0$. To incorporate a mild regularization without significantly slowing down the convergence, we propose to use a small, constant value of $\lambda$ throughout the algorithm. We suggest $\lambda=0.001$ as a practical default value, but to obtain optimal performance the value will need to be tuned on a case-by-case basis; see Fig. 6 for some numerical results for different values of $\lambda$. We refer to the resulting scheme as the regularized block Kaczmarz method, or ReBlocK.
@@
(24) $\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$
@@
(25) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\frac{\left\lVert r^{(\rho)}\right\rVert^{2}}{m}.$
@@
(26) $\left\lVert r^{(\rho)}\right\rVert^{2}\leq\left(1+\frac{N}{\lambda}\right)\left\lVert r\right\rVert^{2}.$
Note that the values of $x^{(\rho)}$ and $\alpha$ here differ from those in Section 2 due to the different choice of $M(A_{S})$ used by ReBlocK. The advantage relative to RBK-U is that ReBlocK-U is able to control both the variance of the iterates and the conditioning of the weight matrix $\overline{W}$ in terms of the reciprocal of the regularization parameter $\lambda$. As a result, the algorithm converges robustly even when the problem contains many nearly singular blocks $A_{S}$. It is worth noting that these advantages could also be attained by truncating the singular values that are smaller than $\lambda$ in the RBK iteration (5). However, the ReBlocK iteration is additionally justified by the fact that it is cheaper to implement than the RBK iteration; see Section 4.1 and Fig. 8 for further discussion of this point.
The proof of Theorem 4.1 is provided in Appendix C. Relative to the proof of Theorem 3.1, the proof of Theorem 4.1 is more complicated because the ReBlocK iteration does not lead to an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$. Instead, the proof relies on a bias-variance decomposition inspired by [12, 23, 18], which also leads to the extra factors of $2$ in the convergence bound.
We are not yet able to analyze the convergence rate parameter $\alpha$ of ReBlocK-U, even in the case of Gaussian data, as to our knowledge there is no existing work bounding the quality of a regularized Gaussian sketch. However, convergence to the ordinary least-squares solution is obtained for a broader class of noisy linear least-squares problems; see Appendix D. Additionally, a fast rate of convergence to the ordinary least-squares solution can be shown when sampling from an appropriate determinantal point process; see Appendix E.
@@ 4.1 Implementation Details
To implement the ReBlocK iterations (7), we directly calculate the $k\times k$ matrix $A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI$ and then use a Cholesky-based linear system solver to calculate $(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x)$. This computation is stable as long as $\lambda$ is not chosen to be too small. Using this approach, the most expensive part of the ReBlocK iteration is calculating $A_{S_{t}}A_{S_{t}}^{\top}$, which has an asymptotic cost of $O(nk^{2})$ just like RBK. Nonetheless, in practice the matrix-matrix multiplication for ReBlocK can be significantly cheaper than the SVD used for RBK. For example, in the experiments of Section 5, ReBlocK iterations are over twenty-five times faster than RBK iterations, as reported in Fig. 8. For the largest problems, the ReBlocK iterations could be further accelerated using iterative solvers; see for example Section 4.1 of [16].
@@ 4.2 Numerical Demonstration
We now test the ReBlocK-U algorithm on two problems whose columns are discretized representations of continuous functions, leading the matrices $A$ to contain many nearly singular blocks of rows. We set $\lambda=0.001$ and generate the inconsistency using random Gaussian noise. We find that TA-RBK-U is unstable for these problems, as expected. Furthermore, TA-ReBlocK-U converges much faster than TA-mSGD in the presence of rapid singular value decay. To understand the impact of the inconsistency on the algorithms, we re-run the case with rapid singular value decay with varying levels of inconsistency, with results in Fig. 4. We find qualitatively similar results in every case. Further details on these experiments can be found in F.1.
@@
Figure 4: Exploration of the effect of the inconsistency on the behavior of the algorithms. The variance of each entry of the noise vector is set to $61e-$ 6 (top left), $41e-$ 4 (top right), and $21e-$ 2 (bottom). The top right panel here is identical to the right panel of Fig. 3, and is reproduced here only for comparison with the top left and bottom.
@@ 5 Natural Gradient Optimization
Our original motivation for this work comes from the problem of training deep neural networks using natural gradient descent [1], which is based on an efficient natural gradient induced by a problem-dependent Riemannian metric. Natural gradient descent has been studied extensively in the machine learning community; see [28, 38, 27]. Furthermore, there is increasing evidence that natural gradient methods can improve the accuracy when training neural networks to solve physical equations. See [29, 9] for applications to physics-informed neural networks and [35, 40] for applications to neural network wavefunctions.
@@
(27) $\min_{\theta}L(\theta),\quad L(\theta)\coloneqq\frac{1}{2}\int_{\Omega}(f_{\theta}(s)-f(s))^{2}{\rm d}s,$
@@
(28) $\theta\leftarrow\theta-\eta G_{N},\quad G_{N}\coloneqq F^{-1}\nabla_{\theta}L(\theta),$
@@
(29) $F=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)\nabla_{\theta}f_{\theta}(s)^{\top}{\rm d}s=J^{\top}J,$
@@
(30) $\nabla_{\theta}L(\theta)=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)(f_{\theta}(s)-f(s)){\rm d}s=J^{\top}[f_{\theta}-f].$
@@
Calculating the natural gradient direction $G_{N}$ using (28) requires a linear solve against the $n\times n$ matrix $F$, which is very challenging in realistic settings when $n\geq 10^{6}.$ This has motivated the development of approximate schemes such as [28]. An alternative approach is to reformulate $G_{N}$ using the structure of $F$ and $\nabla_{\theta}L(\theta)$:
(31) $G_{N}$ $=(J^{\top}J)^{-1}J^{\top}[f_{\theta}-f]$
(32) $=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Jx-[f_{\theta}-f]\right\rVert^{2},$
where the norm in the final expression is the $L_{2}$-norm in function space. This least-squares formulation has been pointed out for example by [27, 7, 19], with the work of Chen and Heyl empowering major advances in the field of neural quantum states. A major goal of the current work is to provide a more solid foundation for the development of natural gradient approximations along these lines.
The natural way to access the data when training a neural network is to sample a set of points in the domain $\Omega$ and evaluate the target function $f$, the network outputs $f_{\theta}$, and the network gradients $\nabla_{\theta}f_{\theta}$ at the sampled points. This is precisely equivalent to sampling a small subset of the rows of (32), which motivates the consideration of row-access least-squares solvers for calculating natural gradient directions. Furthermore, both empirical evidence from scientific applications [33, 45] and theoretical evidence from the literature on neural tangent kernels [4, 39, 5] suggest that $J$ should be expected to exhibit fast singular value decay, motivating the possibility that RBK and ReBlocK could converge rapidly when solving (32). Note that these observations translate straightforwardly from the simple function learning problem (27) to the realistic problems of training physics-informed neural neteworks or neural network wavefunctions.